How many rupees will be the compound interest on Rs 16000 for 9 months at 20% per annum compounded quarterly ?
Answer by addingup(3677)

Using the compound interest formula (this formula adds the interest to your amount but using it means one less formula to memorize)
:
16000(1+0.20/4)^3 =
16000(1.1576) = 18521.60
18521.60-16000 = 2521.60 Rs.
:
Finding the compound interest only:
:
16000[((1+0.20/4)^3)-1]
16000(1.1576-1)
16000*0.1576 = 2521.60
:
Either way works, whatever is easier for you to memorize.
John